If the range of the function
y=(1/5)x+1
is {-2, 3, 6}, then what is the corresponding domain?
A. {-16, 9, 24}
B. {-15, 10, 26}
C. {-11, 14, 29}
D. {-15, 10, 25}

Answer by Edwin McCravy(20056)

y=(1/5)x+1
The range is {-2, 3, 6}, 

Substitute y=-2 in

y=(1/5)x+1
-2=(1/5)x+1
Multiply both sides by 5 to clear of fractions:

-10=x+5
-15=x

So -15 is in the domain.  This rules out A and C

Substitute y=3 in

y=(1/5)x+1
3=(1/5)x+1
Multiply both sides by 5 to clear of fractions:

15=x+5
10=x  so 10 must be in the domain, so it's still either B or D

Substitute y=6 in

y=(1/5)x+1
6=(1/5)x+1
Multiply both sides by 5 to clear of fractions:

30=x+5
25=x

So 25 must be in the domain.

Answer: D. {-15, 10, 25} 

Edwin